OpenStudy (anonymous):

Use a matrix to find the solution to the system of equations. -8x - 8y = -16 6x - 9y = -108

OpenStudy (harsimran_hs4):

\[\left[\begin{matrix}-8 & -8 \\ 6 & -9\end{matrix}\right] \left(\begin{matrix}x \\ y\end{matrix}\right) = \]\[\left(\begin{matrix}-16 \\ 108\end{matrix}\right)\]

OpenStudy (harsimran_hs4):

now you need to find inverse of first matrix or are you willing to involve determinants also

OpenStudy (anonymous):

how do you find the inverse of a matrix?

OpenStudy (harsimran_hs4):

many ways tough if you could tell me which method you have been using so far to solve such questions

OpenStudy (anonymous):

i honestly dont know :/

OpenStudy (anonymous):

i've been sick for 2 weeks and im so far behind i dont know what im doing..

OpenStudy (harsimran_hs4):

can you involve determinants also or no

OpenStudy (anonymous):

like the answer choices are as follows (–6, 8) (6, 8) (8, –6) (6, –8)

OpenStudy (anonymous):

What is a determinant?

OpenStudy (harsimran_hs4):

i know they would be of this form.....because finally it will give x and y

OpenStudy (anonymous):

I'm sorry if im being difficult :/ im honestly just so lost im kind of panicing

OpenStudy (harsimran_hs4):

just take out time and spend around 2 hrs on text i am posting believe me this will clear your concepts and you will be able to solve any such equations

OpenStudy (anonymous):

Thanks so much, lifesaver! haha

OpenStudy (harsimran_hs4):

:)

OpenStudy (harsimran_hs4):

if you are clear with basic matrices concepts you can jump to inverse part in the text but if you have time then i suggest go through it fully

OpenStudy (anonymous):

Alright thanks a lot honestly